OpenStudy (anonymous):

If K = {(x, y)|x - y = 5}, find the corresponding range of y for the domain {0, 2, 4}. {-5, -3, -1} {0, 2, 4} {5, 7, 9}

OpenStudy (anonymous):

iil help

OpenStudy (anonymous):

The domain is given to you, so that's easy: Domain: x = {0, 2, 4} K is a set defined by all values x and y such that x - y = 5. x - y = 5 means -y = 5 - x y = x - 5 If x = 0, y = 0 - 5 = -5 If x = 2, y = 2 - 5 = -3 If x = 4, y = 4 - 5 = -1 So the range is y = {-5, -3, -1 }